OpenStudy (studentchaitrali):

The resultant of two vectors u and v is perpendicular to u and it's magnitude is half of magnitude of vector v, the angle between u and v in degrees is: 120 60 90 150

OpenStudy (shamim):

Can u draw it ?

OpenStudy (shamim):

I wanna solve this problem by parallelogram law of vector addition

OpenStudy (isaidavila):

You have to do it by the sine law V/sin90=(v/2)/sin@ Sin@=1/2 @=30 Since it is a parallelogram opposite angles are equal 90+60= 150 Don't know how can I put an image here so I make the best describing you the solution

OpenStudy (vincent-lyon.fr):

As @shamim said, a drawing is the right way to tackle the problem.